Calculate the following expressions while converting the first argument to the number system of the second one and present the result in decimal number system:

(a) 156 + 1515

(b) 1213 × 1212

(c) 1010 − 10100

(d) 712 × 127


Note: treat ab as number a presented in number system of base b.
Answer #1

a)

156 + 1515 = B15 + 1515 = 2215 = 3210

b)

1213 × 1212 = 1312 * 1212 = 19612 = 25810

NOTE: As per HOMEWORKLIB POLICY, I am allowed to answer only 2 questions (including sub-parts) on a single post. Kindly post the remaining questions separately and I will try to answer them. Sorry for the inconvenience caused.